what is the discriminant of the quadratic equation for 4x^2-9x=0 and what is the number and type of solution for free
Answer by Cabdiqafaar(9)

4x^2-9x=0
use factorisation method
x(4x-9)=0
x=0
4x-9=0
divide both sides by 4
x=9/4